Solve the system of equations by the substitution method
x + 3y = 32
-3x + 2y = 3

Answer by algebrahouse.com(1659)

x + 3y = 32 -------> x = -3y + 32 {subtracted 3y from both sides}
-3x + 2y = 3

-3x + 2y = 3 {top equation}
-3(-3y + 32) + 2y = 3 {substituted -3y + 32, in for x, into top equation}
9y - 96 + 2y = 3 {used distributive property}
11y - 96 = 3 {combined like terms}
11y = 99 {added 96 to both sides}
y = 9 {divided both sides by 11}

x = -3y + 32 {re-arranged top equation}
x = -3(9) + 32 {substituted 9, in for y, into new top equation}
x = -27 + 32 {multiplied -3 by 9}
x = 5 {added}

x = 5 and y = 9

For more help from me, visit: www.algebrahouse.com